a) Two spheres, each with a charge of +3x10-6C, are separated by 100 mm. Are they attracting each other or rejecting? Determine the force between them. 

b) A charge +4x10-6C placed at point P within an electric field, experiences a force of F=-8x10-4N. Determine the intensity of the electric field.




1
Expert's answer
2021-10-20T16:46:35-0400

Part(a)

"E=\frac{F}{q}"

Put value

"E=\frac{-8\times10^{-4}}{4\times10^{-6}}=-200N/C"

Part(b)


"F=\frac{kq_1q_2}{d^2}=\frac{9\times10^9\times3\times10^{-6}\times3\times10^{-6}}{(10^{-2})^2}=810N"